OpenStudy (anonymous):

f(x) = x2 - 16 and g(x) = x+4. Find F over G of and its domain

OpenStudy (anonymous):

\[\frac{ f(x) }{ g(x) }\] is this what you mean?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[\frac{ (x^2 -16) }{ (x+4) }\] you can factor the numerator in a way so as to get two expressions one of which will be (x+4) which will cancel out with the denominator

OpenStudy (anonymous):

so in other words (x+4)(?) = (x^2 - 16)

OpenStudy (anonymous):

4x and 16x^2 right

OpenStudy (anonymous):

im sorry i don't quite understand your expression can you write it out?

OpenStudy (anonymous):

im not getting it

OpenStudy (anonymous):

okay do you know how to factor \[(x^2-16)\]

OpenStudy (welshfella):

x^2 - 16 is the difference of 2 squares x^2 is a perfect square and so is 4

OpenStudy (welshfella):

have you factored something like that before?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

are you familiar with the difference of squares method (its the way to factor expresions like these )

OpenStudy (anonymous):

yeah but its a real number right

OpenStudy (anonymous):

basically it means that if your constant (in this case the 16) is a perfect square you can factor the expression by multiplying two expressions which's constant is the perfect square of your original constant

OpenStudy (anonymous):

okay im over complicating this XD

OpenStudy (anonymous):

okay do you know what \[\sqrt{16} = ?\]

OpenStudy (anonymous):

2*8

OpenStudy (anonymous):

umm no , what number squared equals 16

OpenStudy (anonymous):

256

OpenStudy (anonymous):

no thats 16 squared

OpenStudy (welshfella):

no what number when multiplied by itself equals 16 like 3 * 3 = 9 So 3 is the square root of 9.

OpenStudy (anonymous):

4

OpenStudy (welshfella):

right

OpenStudy (anonymous):

correct which means that \[\sqrt{16} = 4\] and because 4 is a whole number we call 16 a perfect square

OpenStudy (anonymous):

okay thanks